24 November 2009 Lab Meeting
- Oakridge deadline is 1 Feb 2010.
‡ Flu R01:Integration
† HDA (Lead: Jaephil, Team:Sonali)
- Start planning R01 for Submission on 2/5/10. MM, JD, CMK.
- Dry reagent storage for RT and HDA - COP test well array provided fo Sonali
- Set-up meeting for 2nd HDA design - Target (Flu or Bacteria), include RT or not, Heat source, address evaporation loss, Reagent delivery and storage solution. Consider marginating flows and include RT?
- Paper resubmission. Cathie will do Wed.
† Sample Concentration (Lead: Jane, Team: Jaephil)
- New design sent to make Rubber mold: outline octagon problem. Revised design sent yesterday.
- Jane will train Cassidy once more on Wednesday and Thursday this week.
- MM 040808 A/PR/8/34 flu stock with PFU = 2X10^4 available to Jane to train Cassidy in ultracentrifuging all tubes
- of virus, pour out sucrose,pipette remaining with 20uL pipette tip and pool for plaque assay for supplying virus
- Jane working on the cell lysate control. - have many tubes now, done
- Set up COMSOL and StarCD, look for governing equations for simulation of evaporation. - ongoing
† SPE Column Optimization for RNA. (Lead: Sonali, Team: Hussam, Jessie, Brendan)
- pH the sample. Optimum 6.6. Make sure that we are there. 3M GUSCN stands at pH 4.2 very acidic (some NaOH)
- "Measure again 1:1 with DEPC water, 1:1 with PBS"
- Silica nanoparticles from Fluka. 15nm "order"
- Still working on the 150nm particles 1X experiments with lambda. HM
- PLanning net silica per channel experiments. with 700nm, 3x possible. Deliver around 11/5/09.
- 150nm and 700nm silica with 0.5mm (3.6uL SPE) channels.
- 150nm Silica did not improve Ct on average, but one channel with 700nm silica performed exceptionally well, which brought the average up.
- "relationship between PFU/ml and copy # of RNA. Do this via a serial dilution of Virus, put through Qiagen kit. Assuming that Qiagen is 100% efficient."
- "Jessie get trained for real on SEM.*"
† PCR - CMI (Lead: Qingqing)
- QQ will work on the initial integration steps of SPE + RT (reservoir)+ PCR.
- SPE+RT+PCR chip is designed.
- toxin A.
HDA assay for toxin A is tested in tube. It works
HDA chip will be designed, and tested.
SPE DNA from lisa for test?
- toxin B.
on chip PCR works.
- PCR2 Paper formatted for LOAC this week.
need to get back from Cathie
- CMK: PCR 1 draft. Analytical Chem. MCK is running more simulations.
‡C. diff Project (Cathie, Sonali, Satish Singh, Lisa J., His post doc )
- Sonali to train Lisa on SPE.
- They don't have syringe pumps and have not asked for training yet. I offered training here which was declined.
- QQ to run test PCR on chip with genomic DNA and Toxin B primers.
‡ Coulter Flu Fraunhofer Project (Lead: Sonali, Team: Sonali, Jessie, Cathie, CMI Folks, Qingqing)
- Meeting 9am Monday (every other week)
- "ASB will give action items."
‡ Agilent Automated Sample Preparation (Lead: Alex)
- So far tested on SNAP2:
- B. subtilis gDNA
- B. subtilis cells, high Cts for the range tested.
- Planned tests SNAP2:
- B. subtilis cells (higher concentrations)
- MDCK gDNA
- HotDog Sample
- NANO-C: Straw making with two plugs in progress, first lysis test with B.Subtillis cells planned for next week.
- Paper on Yeast: Extended Literature Review (12/11)
- PATH: straw holder for Blackbird is machined
‡ COBRA (Lead: Jaephil, Team: Cathie, Jane for virus only, PHO folks)
- New metal piece - First week of Dec.
- Evap paper 1st draft by the end of this week
‡ Biointerfaces group (Lead: MinCheol, Team: Cathie, MCK, Wong and Meller folks)
- New experiments happening now.
- Cathie will submit paper inquiry.
‡ CIMIT- Sepsis (Lead:Cathie, Team:TBA)
- IRB approved.
- Cathie submitted the companion BU IRB form for exemption. Revised, still still waiting.
‡ PATH Grant (Lead:Cathie, Team:Frank, Sean, Mark, Jake Trueb (ME))
- Phone meeting with PATH this week. 10/13.
- Senior Project Proposals submitted; proposal presentations Friday after Thanksgiving (Dec. 4).
- Frank: 7- and 14-day unstabilized RNA results emailed to group last week.
- Frank: PVA experiments to begin after Thanksgiving.
- Sean: Design has been submitted to Alex and awaiting the product. I can then get back to working on the optimization of the straws, ie the best monolith composition for extracting DNA/RNA onto the straws
- Sean met with Alex for an integration of straws on black machine. They sketched a design. Sean will make a professional drawing and create a required parts list. Alex will supervise the design. Before handing in the job to the shop floor, the design will be presented to Cathie.
- Mark to begin initial DNA testing with shorter time periods (1,3 days, etc.) to be completed this(?) week. PCR training completed. Ready to begin testing. Waiting for new straw fixture to be ready for testing (where will it be stationed?).
‡IIH Senior Project.
- Trouble locating PMMA sheet.
- Try Dow Chemical Website. FOOD safe, FDA grade.